40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2977     Entry time: Tue May 25 00:10:24 2010
Author: Zach 
Type: Update 
Category: IOO 
Subject: IMC table leveled again 

The IMC table had to be leveled again, for two reasons: 1) It was un-leveled when Jenne and Kiwamu removed an extra beam dump when they took the beam profile measurements, and 2) the stack of weights I had put there before was too tall to allow the beam to pass (I didn't realize that the BS chamber is offset a bit to the north, so the beam passes right over the NE edge of the IMC table).

First of all, I was wrong before when I said that the stack of weights was 4 blocks tall; it was 6 blocks tall. I re-leveled the table this afternoon by removing the top three blocks and placing them immediately south of the bottom three in the original stack, while also moving the circular weight north of its previous position. The table is now balanced roughly to within the tolerance of the bubble level I was using.

After the leveling, I tried to re-lock the modecleaner. Upon removing the beam block on the PSL table, I got some sort of resonance flashes on the MC TRANS monitor. With some minor adjustments to MC2&3, I was able to get a decent TEM00 mode to hit. The cavity wouldn't lock, so I went to the AP table and checked to make sure that the REFL beam was hitting the PD. It was, but the beam was very close to the edge of the focusing lens, so I moved the steering mirror slightly to make the situation a little better.

I then went to the control room to finish the by-now-mundane task of fine-tuning the MC lock, but today's was a worthier opponent. For some reason, the thing didn't want to lock for more than a few seconds at a time. I saw that the spot on MC2 was quite a bit off-center, so I ran the MC2_spot_xxx scripts to get it visually in place, then revisited the AP table to ensure that the REFL beam was still on the PD. No dice.

I don't know what was different. I had Ameristat over the opening between the tanks, with posters on top and on the sides (as usual), and I checked to ensure that the servo gains were at the appropriate levels. Joe pointed out that IOO VME was not responding, but we didn't seem to think that this was the problem (based on nothing I can put in words or stick figure cartoons), and the "alive" indicator on the Auto-Lock control in the MEDM screen was not blinking, as it usually is, but I don't know what bearing this has on anything.

I will try to lock again tomorrow.

 

ELOG V3.1.3-